# On canonical differential equations for Calabi-Yau multi-scale Feynman integrals

### v1

@@ 3.3 $l=3$, Calabi-Yau three-fold Feynman integrals
$M_{n+m-1}$ $=\dfrac{1}{\epsilon}\dfrac{\partial}{\partial{\tau_{\alpha}}}M_{h^{2,1}+\beta}+\sum_{k=0}^{n+m-2}F_{(n+m-1),k}\,M_{k}\,,\qquad$ (97e)
@@
The structure of this differential equation is once more derived from the the Griffiths transversality conditions from Eq. (42). This corresponds to the structure we reviewed in Eq. (55).
The differential equation (100) coincides with the differential equation for the unipotent part of the period matrix from section 3.2 in ref. Ducker:2025wfl , showing that effectively taking $\tau_{a}$ derivatives, corrected by the appropriate cubic coupling $C_{a,b,c}$, amounts to multiplying the differential equation matrix by the inverse of the semi-simple part of the Wronskian. The $\epsilon$-rescaling is done similarly as in the one variable case, hence, as reviewed in paragraph 2.6, corresponds to the $\epsilon$-rescaling from refs. Gorges:2023zgv ; Duhr:2025lbz . Finally, the auxiliary functions $F_{i,j}$ are required to cancel the terms not proportional to $\epsilon$, as in refs. Gorges:2023zgv ; Duhr:2025lbz .
@@ 4.1 The democratic approach
Explicit formulas for $F_{5,4}^{\text{dem}}$, $F_{6,4}^{\text{dem}}$ and $F_{7,4}^{\text{dem}}$ as well as the resulting connection matrix $\tilde{\mathcal{A}}$, close to the MUM-point $\mathbf{x}=0$, can be found in the ancillary file sunrise.nb. We call this democratic approach because we take derivatives not only with respect to $\tau$ but also with respect to punctures, placing all the moduli on equal footing. With this basis the differential equation is in canonical form. We give all results in terms of elliptic functions and the functions $F_{i,j}^{\mathrm{dem}}$ are determined leveraging the modular properties of the functions (see section 5.2 of ref. Giroux:2022wav or section 3.1 of ref. Giroux:2024yxu ).
@@ 4.2 The decoupling approach
where
@@
The terms $F_{i,j}^{\text{dec}}$ which we have collected in ancilliary file sunrise.nb will nevertheless introduce the entries of the Jacobian $\mathfrak{J}_{i,j}$ of Eq. (114). We therefore interpret that, the first step of the algorithm (which $\epsilon$-factorises the diagonal terms) solely introduce $\tau$ by holding $x_{2}$ and $x_{3}$ constant. Then, the second step (which $\epsilon$-factorises the remaining off-diagonal part by introducing the functions $F_{i,j}$) will in turn introduce the punctures $z_{1}$ and $z_{2}$. With this basis the differential equation is in canonical form, close to the MUM-point at $\mathbf{x}=0$. Note that we give all results in terms of elliptic functions and the functions $F_{i,j}^{\mathrm{dec}}$ are determined leveraging the modular properties of the functions (see section 5.2 of ref. Giroux:2022wav or section 3.1 of ref. Giroux:2024yxu ).
@@ 5.1 Banana three-loop with two internal unequal masses
however, since our manifold depends only on $\text{dim}(\mathbf{z})=2$ moduli ($\mathbf{z}=(z_{1},z_{2})$), there are ${20-2=18}$ vanishing entries in the period vector. Hence we define
@@ 5.1.1 The 3-1 configuration
00 0 $=\Bigl{(}\frac{\partial^{3}}{\partial z_{1}^{3}}+q^{(0)}_{1,1(z_{1},z_{1},z_{1})}\,\frac{\partial^{2}}{\partial z_{1}^{2}}+q^{(0)}_{1(z_{1},z_{1},z_{1})}\,\frac{\partial}{\partial z_{1}}+q^{(0)}_{2(z_{1},z_{1},z_{1})}\,\frac{\partial}{\partial z_{2}}+q^{(0)}_{(z_{1},z_{1},z_{1})}\Bigr{)}\,I_{1,1,1,1,0,0,0,0,0},$ (131)
00 0 $=\Bigl{(}\frac{\partial^{3}}{\partial z_{2}\partial z_{1}^{2}}+q^{(0)}_{1,1(z_{2},z_{1},z_{1})}\,\frac{\partial^{2}}{\partial z_{1}^{2}}+q^{(0)}_{1(z_{2},z_{1},z_{1})}\,\frac{\partial}{\partial z_{1}}+q^{(0)}_{2(z_{2},z_{1},z_{1})}\,\frac{\partial}{\partial z_{2}}+q^{(0)}_{(z_{2},z_{1},z_{1})}\Bigr{)}\,I_{1,1,1,1,0,0,0,0,0}.$
@@ 5.2 Banana four-loop with two internal unequal masses
$\dfrac{\partial}{\partial\tau_{i}}\dfrac{\partial}{\partial\tau_{j}}\dfrac{\psi_{2}^{(k)}}{\psi_{0}}=C_{i,j,k}\,,\quad\dfrac{\partial}{\partial\tau_{i}}\dfrac{\partial}{\partial\tau_{j}}\dfrac{\psi_{3}}{\psi_{0}}=\tau_{1}\,C_{i,j,1}+\tau_{2}\,C_{i,j,2}\,.$ (154)
@@
=$=\biggl{\{}\delta_{j,1}+\delta_{j,2}\tau_{1}+\delta_{j,3}\tau_{2}+\delta_{j,4}\dfrac{\partial}{\partial\tau_{1}}\mathcal{F}+\delta_{j,5}\dfrac{\partial}{\partial\tau_{2}}\mathcal{F}$

### v2

@@ 3.3 $l=3$, Calabi-Yau three-fold Feynman integrals
$M_{n+m-1}$ $=\dfrac{1}{\epsilon}\dfrac{\partial}{\partial{\tau_{\alpha}}}M_{h^{2,1}+\beta}+\sum_{k=0}^{n+m-2}F_{(n+m-1),k}\,M_{k}\,$$\qquad$ (97e)
@@
The structure of this differential equation is once more derived from the Griffiths transversality conditions from Eq. (42). This corresponds to the structure we reviewed in Eq. (55).
The differential equation (100) coincides with the differential equation for the unipotent part of the period matrix from section 3.2 in ref. Ducker:2025wfl, showing that effectively taking $\tau_{a}$ derivatives, corrected by the appropriate cubic coupling $C_{a,b,c}$, amounts to multiplying the differential equation matrix by the inverse of the semi-simple part of the Wronskian. Finally, the $\epsilon$-rescaling and the introduction of the auxiliary functions $F_{i,j}$ is done similarly as in section 3.2.
@@ 3.4 $l>3$, Calabi-Yau $l$-fold Feynman integrals
Note that the auxiliary functions $F_{i,j}$ introduced throughout this section, are defined by a set of differential equations and can, in general, have a complicated analytic structure that can go beyond periods (and their derivatives) of Calabi-Yau manifolds paperDM. For the purpose of this work, these differential equations will (except for the elliptic case) only be solved as a series expansion around the MUM-point. We leave for future work Duhr:2025xyy, a more in-depth discussion of these auxiliary functions.
@@ 4.1 The democratic approach
Explicit formulas for $F_{5,4}^{\text{dem}}$, $F_{6,4}^{\text{dem}}$ and $F_{7,4}^{\text{dem}}$ as well as the resulting connection matrix $\tilde{\mathcal{A}}$, close to the MUM-point $\mathbf{x}=0$, can be found in the ancillary file sunrise.nb. We call this democratic approach because we take derivatives not only with respect to $\tau$ but also with respect to punctures, placing all the moduli on equal footing. With this basis the differential equation is in canonical form. We give all results in terms of elliptic functions and the functions $F_{i,j}^{\mathrm{dem}}$ are determined leveraging the modular properties of the functions as in Refs. Giroux:2022wav; Giroux:2024yxu. We briefly review this method in appendix C.
@@ 4.2 The decoupling approach
where [^15] Note that the choice of variable with respect to which we take a derivative is arbitrary as long as the periods (105) depend on it.
@@
The terms $F_{i,j}^{\text{dec}}$ which we have collected in ancilliary file sunrise.nb will nevertheless introduce the entries of the Jacobian $\mathfrak{J}_{i,j}$ of Eq. (114). We therefore interpret that, the first step of the algorithm (which $\epsilon$-factorises the diagonal terms) solely introduce $\tau$ by holding $x_{2}$ and $x_{3}$ constant. Then, the second step (which $\epsilon$-factorises the remaining off-diagonal part by introducing the functions $F_{i,j}$) will in turn introduce the punctures $z_{1}$ and $z_{2}$. With this basis the differential equation is in canonical form, close to the MUM-point at $\mathbf{x}=0$. The functions $F_{i,j}^{\mathrm{dec}}$ are determined as in section 4.2.
@@ 4.3 Comparing the results
We therefore find that in this given case, there is a freedom on the starting set of master integral we can use. It turns out that here, both sets of master integrals are easy to determine from the integral reduction. The main difference between these choices comes from the $\epsilon$-factorisation procedure as the decoupling approach shifts all the functions depending on the marked points into the functions $F_{i,j}^{\text{dec}}$.
The democratic approach, on the other hand, already incorporates the marked points of Eq. (108) the derivatives with respect to the moduli. This is also evidenced by the ansatz of Eq. (4.2) having 8 functions $F_{i,j}^{\text{dec}}$ as opposed to the 3 functions $F_{i,j}^{\text{dem}}$ of Eq. (4.1). It is therefore a question of preference whether one prefers to make the marked points manifest as in the democratic approach, or whether one keeps this information in the functions $F_{i,j}^{\text{dec}}$.
@@ 5.1 Banana three-loop with two internal unequal masses
however, since our manifold depends only on $\text{dim}(\mathbf{z})=2$ moduli ($\mathbf{z}=(z_{1},z_{2})$), there are 202=18{20-2=18} vanishing entries in the period vector. Hence we define
@@ 5.1.1 The 3-1 configuration
$0$ $=\Bigl(\frac{\partial^{3}}{\partial z_{1}^{3}}+q^{(0)}_{1,1(z_{1},z_{1},z_{1})}\,\frac{\partial^{2}}{\partial z_{1}^{2}}+q^{(0)}_{1(z_{1},z_{1},z_{1})}\,\frac{\partial}{\partial z_{1}}+q^{(0)}_{2(z_{1},z_{1},z_{1})}\,\frac{\partial}{\partial z_{2}}+q^{(0)}_{(z_{1},z_{1},z_{1})}\Bigr)\,I_{1,1,1,1,0,0,0,0,0},$ (131)
$0$ $=\Bigl(\frac{\partial^{3}}{\partial z_{2}\partial z_{1}^{2}}+q^{(0)}_{1,1(z_{2},z_{1},z_{1})}\,\frac{\partial^{2}}{\partial z_{1}^{2}}+q^{(0)}_{1(z_{2},z_{1},z_{1})}\,\frac{\partial}{\partial z_{1}}+q^{(0)}_{2(z_{2},z_{1},z_{1})}\,\frac{\partial}{\partial z_{2}}+q^{(0)}_{(z_{2},z_{1},z_{1})}\Bigr)\,I_{1,1,1,1,0,0,0,0,0}.$
@@ 5.2 Banana four-loop with two internal unequal masses
$\dfrac{\partial}{\partial\tau_{i}}\dfrac{\partial}{\partial\tau_{j}}\dfrac{\psi_{2}^{(k)}}{\psi_{0}}=-C_{i,j,k}\,,\quad\dfrac{\partial}{\partial\tau_{i}}\dfrac{\partial}{\partial\tau_{j}}\dfrac{\psi_{3}}{\psi_{0}}=\tau_{1}\,C_{i,j,1}+\tau_{2}\,C_{i,j,2}\,.$ (154)
@@
$=\biggl\{\delta_{j,1}+\delta_{j,2}\tau_{1}+\delta_{j,3}\tau_{2}+\delta_{j,4}\dfrac{\partial}{\partial\tau_{1}}\mathcal{F}+\delta_{j,5}\dfrac{\partial}{\partial\tau_{2}}\mathcal{F}$
@@ Appendix B Explicit proof for K3 Feynman integrals
Appendix C Modular bootstrap
Here we briefly review a method to solve linear differential equations involving modular forms. This method was first introduced in Ref. Giroux:2022wav and applied again in Ref. Giroux:2024yxu. In Eq. (4.1) and Eq. (4.2) we have a set of functions $F_{i,j}$ which $\epsilon$-factorise the differential equation. They follow a linear differential equation and from the elliptic nature of the sunrise integral we know that they are modular forms. To determine them, we will use their modular properties instead of solving the differential equation directly.
When $F_{i,j}$ are modular forms of weight one depending on $r$ variables, we find:
$\text{d}F_{i,j}=R_{i,j}^{(0)}\psi_{0}+\sum_{n=1}^{r}R_{i,j}^{(k)}\dfrac{\partial}{\partial x_{k}}\psi_{0}\,$, (228)
where $R_{i,j}^{(k)}$ are modular invariant one-forms. Since $\psi_{0}$ transforms as a modular form of weight one
$\gamma(\psi_{0})=(c\tau+d)\psi_{0}\,$, (229)
with $\gamma(\tau)=\frac{a\tau+b}{c\tau+d}$ for adbc=1ad-bc=1 being the standard modular transformation, we find that Eq. (228) transforms as
$\text{d}\biggl((c\tau+d)F_{i,j}\biggr)=R_{i,j}^{(0)}(c\tau+d)\psi_{0}+\sum_{n=1}^{r}R_{i,j}^{(n)}\dfrac{\partial}{\partial x_{n}}(c\tau+d)\psi_{0}\,$, (230)
which leads to
$F_{i,j}\,\text{d}\tau=\psi_{0}\sum_{n=1}^{r}R_{i,j}^{(n)}\dfrac{\partial\tau}{\partial x_{n}}\,$, (231)
where the derivatives can be expressed through the Jacobian as $\frac{\partial\tau}{\partial x_{n}}=(\mathfrak{J}^{-1})_{1,n}\,$.
When the functions $F_{i,j}$ are modular forms of weight two we find a differential equation of the form
$\text{d}F_{i,j}=\psi_{0}\biggl(R_{i,j}^{(0)}\psi_{0}+\sum_{n=1}^{r}R_{i,j}^{(k)}\dfrac{\partial}{\partial x_{k}}\psi_{0}\biggr)\,$, (232)
which leads to
$F_{i,j}\,\text{d}\tau=\dfrac{\psi_{0}^{2}}{2}\sum_{n=1}^{r}R_{i,j}^{(n)}\dfrac{\partial\tau}{\partial x_{n}}\,$, (233)
through analogous arguments.